OpenStudy (anonymous):

x^3+3x^2+25x+75 zeros:-5i, 5i What is the remaining zero? I looked up how to do this and it said to try long division and I did that and got a remainder of 25. Now I'm confused on what to do next

OpenStudy (anonymous):

Where does the x-3 come from?

OpenStudy (anonymous):

Yeah I saw that after that I tried all these other ways. I feel so dumb lol but thank you

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

Divide the polynomial by ( x+5i)(x-5i)= x^2+25 You find the ratio to be (x+3) so the other root is x =-3